

********************************************************************
UARS WEEKLY STATUS REPORT - 05 NOV 2004
********************************************************************
***********
SOLSTICE
***********
This week, SOLSTICE was in daylight only operations from Monday (306)
to mid-day on Thursday (309) in support of the yaw maneuver. Otherwise,
the instrument continues to perform exceptionally, taking solar data on
all available orbital tracking periods. An average of 60% of data were
returned this week accounting for 397 minutes of solar observing time
per day.
***********
SUSIM
***********
SUSIM was powered down Nov. 1-3 to save power during the UARS
yaw-around. Otherwise, the experiment successfully gathered daily
mid (1.1 nm) resolution, MgII, occultation, solar pointing offset,
and continuous monitoring scans for time periods with UARS
telemetry or tape recording.
***********
PEM
***********
HEPSA data files are version 2. Currently, HEPSA data between 1991274
and 2004189 has been reprocessed and archived at the GSFC DAAC. HEPSA data
files from 2004190 to 2004199 are processed and waiting for archiving.
MEPS data files are version 1. Processed MEPS data files at the
DAAC archive are between 1991274 and 2004179. MEPS data files from 2004180
to 2004199 are processed and waiting for archiving.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron data for
the October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality. Proton data from these storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
Requests have been received for storm spectral data. Overview plots were
produced and these plots may be found on the web at
http://climatology.space.swri.edu/UARS_images_postedJune2004.html
for the storms of October 29 of 1991, November 8-9 of 1991, April 6-7 of 2000,
March 31 - April 1 of 2001, April 11-12 of 2001, November 6-7 of 2001, and
November 11-12 of 2001.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 23 October 2004 (Orbit 71782, GMT Day 297) through
29 October 2004 (Orbit 71886, GMT Day 303).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 303/2315 71886 POWER OFF
The beta angle ranged between the angles of 41.8 degrees to 14.5
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 0 degrees on 02 November 2004 (DOY 307).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 0.63 to 0.65 degrees C. Additional
battery performance data for each battery on 22 October and 29 October
2004 are provided in Appendix D.
The clock error ranged from 10.6 to -10.1 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 02
HALOE = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 08
Seven (7) Anomaly Reports remained OPEN at the end of this report
period (see Appendix E).
GROUND SYSTEM OPERATION
Eight (8) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
This week the FOT took their first SMA pass with the real spacecraft.
Even though the CMS is not able to receive the SMA contacts, the FOT
was able to schedule an SMA event on the UPS. The test was successful
with commands and dumps performing properly. Further use of the TDRS
SMA service will be postponed until the transfer issue between UPS and
CMS can be resolved.
The few test SMA events, which were scheduled via the workaround
feature on the UPS, were confirmed by scheduling this week. The test
events were deleted by the FOT, but the test verified the successful
execution of the workaround throughout the scheduling process.
Therefore no modifications to the FDF-provided UAV file will be
necessary.
MEETINGS
The FOT attended a meeting with FDF/Software Development on 21 October
2004 to discuss the rehost of FORMS 3 & 4 and the SMA issues. The
rehost plan was given the go ahead, but action items were taken by
FDF, sustaining engineering and software development to resolve the
file transfer issue between the UPS and CMS. A follow up meeting is
scheduled for 04 November 2004.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 02 November 2004 (DOY 307).
This maneuver will be a Reverse to Forward Flight maneuver (Southern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 29 October 2004
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
71865 302/1336 04-307 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
71884 303/1944 04-308 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
71886 303/2315 04-309 HALOE POWER OFF (YAW)
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 29 October 2004
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 29 October 2004
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
71865 302/1343 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
71884 303/2000 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 29 October 2004
22 October 2004 - GMT Day 296 Beta = 42.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.05/-4.31 +00.0/+00.0 07.0 0.00 0.0
2 11.2/- 50.4 +6.11/+5.14 +25.6/-15.6 26.2 1.04 19.6
3 -11.2/-168.0 +4.50/+3.87 +24.0/-15.2 26.2 1.03 18.8
29 October 2004 - GMT Day 303 Beta = 14.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.05/-4.31 +00.0/+00.0 07.0 0.00 0.0
2 05.6/- 39.2 +5.79/+5.14 +28.4/-13.6 26.2 1.03 20.3
3 -56.0/- 56.0 +4.50/+3.87 +27.2/-13.6 26.2 1.02 19.6
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period
APPENDIX E
UARS ANOMALY REPORTS
Week ending 29 October 2004
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-076 FOT 71783 297/0234 37341 Late Acquisition
04-066 FOT 71809 298/2040 30832 Generic Late Acquisition
04-077 FOT 71811 298/2220 37357 Multiple Data Dropouts
04-078 FOT 71826 299/2326 37358 DATA DROPOUTS
04-079 FOT 71834 300/1203 37396 TDRSS 275 Data Dropouts
04-066 FOT 71836 300/1525 30832 Generic Late Acquisition
04-080 FOT 71849 301/1200 37411 TDRSS 275 Dropouts
04-082 WSGT 71861 302/0647 37419 Drop Outs on the Return Link
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-073 FOT 71691 291/0026 37270 Data Drop Outs
04-074 FOT 71730 293/1330 37291 30 sec Data Dropout
04-075 FOT 71776 296/1549 N/A Canberra Antenna Problem
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-066 FOT See List Oct 04 30832 Generic Late Acquisition
04-076 FOT 71783 297/0234 37341 Late Acquisition
04-077 FOT 71811 298/2220 37357 Multiple Data Dropouts
04-078 FOT 71826 299/2326 37358 DATA DROPOUTS
04-079 FOT 71834 300/1203 37396 TDRSS 275 Data Dropouts
04-080 FOT 71849 301/1200 37411 TDRSS 275 Dropouts
04-082 WSGT 71861 302/0647 37419 Drop Outs on the Return Link